OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3).

OpenStudy (anonymous):

y = 2x - 7 y = - 1/2x - 2 y = 2x - 2 y = -1/2 x - 7

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

@prettygirlswag<3

OpenStudy (anonymous):

@justicehunter

OpenStudy (anonymous):

@_diamondlife_

OpenStudy (anonymous):

@_DiamondLife_

OpenStudy (anonymous):

@PrettyGirlSwag<3

OpenStudy (anonymous):

helppppppppp

OpenStudy (anonymous):

y = 2x + 2 m = 2 solving its y-intercept y = mx + b 3 = 2(5) + b 3 = 10 + b 3 - 10 = b b = - 7 the equation of the line is y = 2x - 7 answer

OpenStudy (anonymous):

omg thank you